What is the area of a triangle in which b = (x + 1)ft and h = 8x ft?
Answer by oberobic(2304)

The area of a triangle is 1/2*base*height.
.
b = base = (x+1) ft
.
h = height = 8x ft
.
Area = 1/2*(x+1)*8x ft^2
.
A = 1/2(8x^2+8x) ft^2
.
A = 4x^2 + 4x ft^2
.
A = 4x(x+4) ft^2
.
Done